OpenStudy (anonymous):

Solve using square roots. Check your answers. 25 x^2 - 1 = 0

OpenStudy (campbell_st):

\[25x^2 = 1\] \[x^2 = 1/25\] \[x = \pm \sqrt{1/25 } = \pm 1/5\]

OpenStudy (ash2326):

We have \[25 x^2-1=0\] Add 1 to both the sides \[25 x^2=1\] Now divide both sides by 25 \[x^2=\frac{1}{25}\]Now take root both sides, we get \[x=\pm \frac{1}{5}\]